2 NaOH + H2SO4 → Na2SO4 + 2 H2O
Working in the lab you spill 20.0 mL of (1.50 moles/1L) sulfuric acid (H2SO4) on the lab bench. You have been warned that sulfuric acid is not safe to touch and you know that you need to clean it up. You teacher informs you that sodium hydroxide will neutralize the spill if you use the correct amount. (1.75 moles/1L) NaOH is available in the lab. Using the chemical equation provided, determine the volume of sodium hydroxide necessary to neutralize the spill so that you can clean up.
1
Expert's answer
2020-05-16T13:45:46-0400

2 NaOH + H2SO4 → Na2SO4 + 2 H2O 

1) m=20.0*1.50*98/1000=2.94 g

2) 2.94*80/98=2.4 g

3) V=m*1000/Mr*C=2.4*1000/40*1.75=34.28 mL NaOH